# Insert-Only versus Insert-Delete in Dynamic Query Evaluation

### v3

@@ 8. Conclusion
This paper puts forward two-way reductions between the dynamic and static query evaluation problems, allowing the transfer of both algorithms and lower bounds between them. For the dynamic problem, it also characterizes the complexity gap between the insert-only and the insert-delete settings. The proposed algorithms recover best known results and produce new ones. Unlike existing lower bounds, the lower bounds obtained in the paper are more general as they do not rely on existing complexity conjectures.
@@ Acknowledgements.
Appendix A Comparison to Prior IVM Approches
@@
  • F-IVM (Nikolic et al., 2020; Kara et al., 2023b): This is a higher-order IVM approach that speeds up the delta computation using a view tree whose structure is modelled on a variable order. Its update times are shown to be at least as good as the prior higher-order IVM approach DBToaster (Koch et al., 2014).
  • IVMϵ (Kara et al., 2020a, 2019): This is an adaptive IVM methodology that uses delta computation and materialized views and also takes the degrees of data values into account. It is adaptive in the sense that it treats updates referring to values with high degree differently from those with low degree. The specific degrees to partition and the partitioning threshold that lead to the best update time can vary depending on the query structure. So far, there is no algorithm that produces the optimal partitioning strategy for arbitrary queries. Therefore, IVMϵ is not a concrete algorithm for arbitrary queries but rather a methodology. It was shown in (Kara et al., 2020a, 2019) how to manually tailor this methodology to the triangle query and several non-hierarchical queries to achieve the worst-case optimal update time.
  • CROWN (Change pROpagation Without joiNs(Wang et al., 2023): This is a higher-order IVM approach tailored to free-connex $\alpha$-acyclic queries. Its maintenance approach is in line with F-IVM, when applied to free-connex acyclic queries, yet it provides a more refined complexity analysis that takes the structure of the update sequence into account. For certain update sequences, such as insert-only or first-in-first-out, the amortized update time is constant. For the purpose of comparing against the other IVM approaches, which do not consider such a fine-grained analysis, we consider the complexity results of CROWN for arbitrary update sequences.
@@
Figure 3 summarizes the update times of all aforementioned IVM approaches. Since CROWN is designed for free-connex acyclic queries, we skip its complexity for the acyclic ones, which are marked “NA”. It is not clear how to manually tailor the IVMϵ approach to all queries, thus some entries are marked “Open”.
@@
$Q_{2p}^{\prime}$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ Open $\mathcal{O}(|\mathcal{D}|^{1/2})$
$Q_{3p}$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|^{1/2})$ $\mathcal{O}(|\mathcal{D}|^{1/2})$
$Q_{\triangle}$ $\mathcal{O}(|\mathcal{D}|^{3/2})$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ NA $\mathcal{O}(|\mathcal{D}|^{1/2})$ $\mathcal{O}(|\mathcal{D}|^{1/2})$
$Q_{\triangle}^{\prime}$ $\mathcal{O}(|\mathcal{D}|^{3/2})$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ NA $\mathcal{O}(|\mathcal{D}|^{2/3})$ $\mathcal{O}(|\mathcal{D}|^{2/3})$
$Q_{\ast}$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ Open $\mathcal{O}(|\mathcal{D}|^{2/3})$
$Q_{3u}$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ Open $\mathcal{O}(|\mathcal{D}|^{2/3})$
$Q_{4u}$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ Open $\mathcal{O}(|\mathcal{D}|^{3/4})$
$Q_{2p}^{\prime\prime}$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|^{1/2})$ $\mathcal{O}(|\mathcal{D}|)$
$Q_{\Box}$ $\mathcal{O}(|\mathcal{D}|^{2})$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ NA $\mathcal{O}(|\mathcal{D}|^{2/3})$ $\mathcal{O}(|\mathcal{D}|)$
Figure 3. Update times of different IVM approaches for the problem $\text{IVM}^{\pm}$. Queries are depicted in Figure 2 and are given by Eq. (14)-(22). MVIVM is our approach, introduced in this paper. CROWN is not applicable to cyclic queries, thus some entries are marked “NA”. IVMϵ is not a concrete algorithm but rather a general methodology. It is not clear how to manually set it up for every query, thus some entries are marked “Open”.
@@
In the following, we illustrate how the aforementioned IVM approaches maintain the 3-path query $Q_{3p}$ and the triangle query $Q_{\triangle}$. We first introduce some useful conventions that simplify the following presentation.
Conventions
In the following, we denote a single tuple update to some relation $R(A,B)$ by $\delta R(a,b)$, which expresses that the tuple $(a,b)$ is inserted to or deleted from $R$. Whenever it simplifies the explanation of an update procedure, we do not distinguish between an insert and a delete and describe the main computation steps common to both cases.
For any materialized relation or view $V(\bm{X})$ with schema $\bm{X}$, we assume that we have indices that allow us to perform the following operations: (1) enumerate the tuples in $V$ with constant delay; (2) check if $V$ contains a given tuple in constant time; (3) insert or delete a tuple in constant time. Similarly, given $\bm{Y}\subseteq\bm{X}$ and a tuple $\bm{y}$ over $\bm{Y}$, we assume that we can: (1) enumerate the tuples in $\pi_{\bm{X}\setminus\bm{Y}}\sigma_{\bm{Y}=\bm{y}}V$ with constant delay; (2) check if $V$ contains a tuple $\bm{x}$ with $\pi_{\bm{Y}}\bm{x}=\bm{y}$ in constant time.
@@ A.4. IVMϵ
Updates can change the degrees of values, which means that a light value can become heavy and vice-versa. The original IVMϵ algorithm moves tuples between the heavy and the light part of a relation whenever the degrees of the values in the tuples pass certain thresholds (Kara et al., 2020a). This leads to additional time cost that can be amortized over update sequences such that the amortized update time remains $\mathcal{O}(|\mathcal{D}|^{1/2})$. Using standard de-amortization techniques, the original IVMϵ algorithm can be tweaked so that the update time is $\mathcal{O}(|\mathcal{D}|^{1/2})$ in the worst case (Kara et al., 2020a).
@@ A.6. MVIVM
Finally, we highlight some of the key insights of our approach, MVIVM, that enables it to maintain the triangle query $Q_{\triangle}$ in time $|\mathcal{D}|^{1/2}$. Example 7.1 presents the high-level idea while Appendix E.1 explains the technical subtleties that arise. At a high-level, our approach maintains $Q_{\triangle}$ by constructing and maintaining the six multivariate extension components $\widehat{Q}_{123},\ldots,\widehat{Q}_{321}$ given by Eq. (4) and (38). Each of these component queries is maintained using Theorem 3.1 along with a variety of technical lemmas described in Appendix E.1. It can be mechanically verified that each of these components has a fractional hypertree width of $\frac{3}{2}$, leading to our amortized update time of $\widetilde{\mathcal{O}}(|\mathcal{D}|^{1/2})$. Let’s take the first component $\widehat{Q}_{123}$ as an example:
@@
Recall that $Z_{1}$, $Z_{2}$, and $Z_{3}$ are temporal attributes representing the lifespans of tuples in $R$, $S$, and $T$. In particular, the lifespan of a tuple in $T(A,C)$ is represented by the concatenation of $Z_{1}\circ Z_{2}\circ Z_{3}$; see Example 6.4. The above tree decomposition corresponds to a query plan where we are projecting $Z_{3}$ out of the child bag before joining with the root bag. Projecting $Z_{3}$ out of the child bag corresponds to “expanding” the lifespans of tuples of $T(A,C)$ from the lifespan $Z_{1}\circ Z_{2}\circ Z_{3}$ to the longer lifespan $Z_{1}\circ Z_{2}$. What this means algorithmically is that we are lazily processing deletes from $T$ in order to avoid doing excessive work for each delete, especially in the case where the same tuple from $T$ is inserted and deleted many times. In particular, a sequence of repeated insert-delete of the same tuple of $T$ corresponds to lifespans whose encodings may share the same prefix $Z_{1}\circ Z_{2}$. Hence, we don’t need to propagate changes from the child bag to the root bag for every such insert-delete because the lifespan $Z_{1}\circ Z_{2}$ remains in the child bag throughout all these insert-deletes. Without this lazy processing, each such repeated insert/delete into $T$ would have required linear update time, thus we would have failed to meet our target update time of $|\mathcal{D}|^{1/2}$. This implicit lazy processing is just one example of sophisticated behavior that comes automatically out of our approach.
@@ Appendix B Missing Details in Preliminaries
Acyclic queries. A query is ($\alpha$-)acyclic if we can construct a tree, called join tree, such that the nodes of the tree are the atoms of the query (coverage) and for each variable, it holds: if the variable appears in two atoms, then it appears in all atoms on the path connecting the two atoms (connectivity(Brault-Baron, 2016).
@@
Width Measures
[…]
Every feasible solution $(\lambda_{R(\bm{X})})_{R(\bm{X})\in\text{\sf at}(Q)}$ to the above program is called a fractional edge cover of $Q$ with respect to $\mathcal{D}$. The optimal objective value of the program is called the fractional edge cover number $\rho^{*}(Q,\mathcal{D})$ of $Q$ with respect to $\mathcal{D}$. The fractional edge cover number gives a bound on the size of $Q(\mathcal{D})$: $|Q(\mathcal{D})|\leq 2^{\rho^{*}(Q,\mathcal{D})}$ (Atserias et al., 2013). This bound is known as the AGM-bound and denoted as $\text{\sf AGM}(Q,D)$. We denote by $\rho^{\ast}(Q)$ the optimal objective value of the linear program that results from the above program by replacing the objective function with $\sum_{R(\bm{X})\in\,\text{\sf at}(Q)}\lambda_{R(\bm{X})}$.
Definition B.2 (Tree Decomposition).
  1. (1)
  2. (2)
Definition B.3 (The restriction $Q_{\bm{Y}}$ of a query $Q$).
Definition B.4 (Fractional Hypertree Width).
$\mathsf{w}(\mathcal{T},\chi)$ =$\stackrel{{\scriptstyle\text{def}}}{{=}}\max_{t\in V(\mathcal{T})}\rho^{*}(Q_{\chi(t)})$
(24) $\mathsf{w}(Q)$ =$\stackrel{{\scriptstyle\text{def}}}{{=}}\min_{(\mathcal{T},\chi)\in\text{\sf TD}(Q)}\mathsf{w}(\mathcal{T},\chi)$
We call a tree decomposition $(\mathcal{T},\chi)$ of $Q$ optimal if $\mathsf{w}(\mathcal{T},\chi)=\mathsf{w}(Q)$. The fractional hypertree width of a union $Q=\bigcup_{i\in[n]}Q_{i}$ of join queries is
$\mathsf{w}(Q)$ =$\stackrel{{\scriptstyle\text{def}}}{{=}}\max_{i\in[n]}\mathsf{w}(Q_{i})$
@@ Appendix C Missing Details in Section 5
In this section, we explain in detail our results on IVM in the insert-only setting, where no deletes are allowed. The main goal is to prove Theorem 3.1 (as well as Theorem 3.3). The results in this section also serve as building blocks for our results on the insert-delete setting in Section E.
@@ C.1. Preliminaries: The Query Decomposition Lemma
Lemma C.1 (Query Decomposition Lemma (Ngo et al., 2014; Ngo, 2018)).
@@
(26) $\sum_{\bm{y}\in\text{\sf Dom}(\bm{Y})}\underbrace{\prod_{R(\bm{X})\in\text{\sf at}(Q)}|R(\bm{X})\ltimes\bm{y}|^{\lambda_{R(\bm{X})}}}_{\text{AGM-bound of $Q\Join\bm{y}$}}\leq\underbrace{\prod_{R(\bm{X})\in\text{\sf at}(Q)}|R|^{\lambda_{R(\bm{X})}}}_{\text{AGM-bound of $Q$}}$

### v4

@@ 8. Conclusion
This paper puts forward two-way reductions between the dynamic and static query evaluation problems that allow us to transfer a wide class of algorithms and lower bounds between the two problems. For the dynamic problem, the paper characterizes the complexity gap between the insert-only and the insert-delete settings. The proposed algorithms recover best known amortized update times and produce new ones. The matching lower bounds justify why our upper bounds are natural. The lower bound function $\omega(Q)$ that we define can be used as a general tool to assess the optimality of some algorithms in database theory.
@@
Although the proposed dynamic algorithms MVIVM are designed to maintain base relations and the query output as sets, they can be extended to maintain bags. As in DBToaster (Koch et al., 2014) and F-IVM (Kara et al., 2024), MVIVM can be extended to maintain the multiplicity of each tuple in a base relation or view and enumerate each tuple in the query output together with its multiplicity.
@@ Acknowledgements.
Appendix A Comparison to Prior IVM Approaches
@@
  • F-IVM (Nikolic et al., 2020; Kara et al., 2024): This is a higher-order IVM approach that speeds up the delta computation using a view tree whose structure is modelled on a variable order. Its update times are shown to be at least as good as the prior higher-order IVM approach DBToaster (Koch et al., 2014).
  • IVMϵ (Kara et al., 2020a, 2019): This is an adaptive IVM methodology that uses delta computation and materialized views and also takes the degrees of data values into account. It is adaptive in the sense that it treats updates referring to values with high degree differently from those with low degree. The specific attributes whose values are partitioned with regard to their degrees and the partitioning threshold that lead to the best update time can vary depending on the query structure. So far, there is no algorithm that produces the optimal partitioning strategy for arbitrary queries. Therefore, IVMϵ is not a concrete algorithm for arbitrary queries but rather a methodology. It was shown in (Kara et al., 2020a, 2019) how to manually tailor this methodology to the triangle query and several non-hierarchical queries to achieve the worst-case optimal update time.
  • CROWN (Change pROpagation Without joiNs(Wang et al., 2023): This is a higher-order IVM approach tailored to free-connex ($\alpha$-)acyclic queries. Its maintenance approach is in line with F-IVM, when applied to free-connex acyclic queries, yet it provides a more refined complexity analysis that takes the structure of the update sequence into account. For certain update sequences, such as insert-only or first-in-first-out, the amortized update time is constant. For the purpose of comparing against the other IVM approaches, which do not consider such a fine-grained analysis, we consider the complexity results of CROWN for arbitrary update sequences.
@@
Figure 4 summarizes the update times of all aforementioned IVM approaches, give in data complexity. The update times of Naïve, Delta, and F-IVM are worst-case and those of CROWN, IVMϵ, and MVIVM are amortized. The update times of MVIVM have an extra $\text{\sf polylog}(|\mathcal{D}|)$ factor, which is hidden by the $\widetilde{\mathcal{O}}$-notation. Since CROWN is designed for free-connex acyclic queries, we skip its complexity for the acyclic ones, which are marked “NA”. It is not clear how to manually tailor the IVMϵ approach to all queries, thus some entries are marked “Open”. In terms of query complexity, MVIVM has an extra factor of $k!$, where $k$ is the number of input relations, due to the use of multivariate extensions (Definiton 3.3).
@@
$Q_{2p}^{\prime}$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ Open $\widetilde{\mathcal{O}}(|\mathcal{D}|^{1/2})$
$Q_{3p}$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|^{1/2})$ $\widetilde{\mathcal{O}}(|\mathcal{D}|^{1/2})$
$Q_{\triangle}$ $\mathcal{O}(|\mathcal{D}|^{3/2})$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ NA $\mathcal{O}(|\mathcal{D}|^{1/2})$ $\widetilde{\mathcal{O}}(|\mathcal{D}|^{1/2})$
$Q_{\triangle}^{\prime}$ $\mathcal{O}(|\mathcal{D}|^{3/2})$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ NA $\mathcal{O}(|\mathcal{D}|^{2/3})$ $\widetilde{\mathcal{O}}(|\mathcal{D}|^{2/3})$
$Q_{\ast}$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ Open $\widetilde{\mathcal{O}}(|\mathcal{D}|^{2/3})$
$Q_{3u}$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ Open $\widetilde{\mathcal{O}}(|\mathcal{D}|^{2/3})$
$Q_{4u}$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ Open $\widetilde{\mathcal{O}}(|\mathcal{D}|^{3/4})$
$Q_{2p}^{\prime\prime}$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|^{1/2})$ $\widetilde{\mathcal{O}}(|\mathcal{D}|)$
$Q_{\Box}$ $\mathcal{O}(|\mathcal{D}|^{2})$ $\mathcal{O}(|\mathcal{D}|)$ $\mathcal{O}(|\mathcal{D}|)$ NA $\mathcal{O}(|\mathcal{D}|^{2/3})$ $\widetilde{\mathcal{O}}(|\mathcal{D}|)$
worst-caseamortized
Figure 4. Update times of different IVM approaches for the problem $\text{IVM}^{\pm}$, give in data complexity. The update times of the approaches Naïve, Delta, and F-IVM are worst-case, while those of CROWN, IVMϵ, and MVIVM are amortized. $\widetilde{\mathcal{O}}$ hides a $\text{\sf polylog}(|\mathcal{D}|)$ factor. Queries are depicted in Figure 3 and are given by Eq. (17)-(25). MVIVM is our approach, introduced in this paper. CROWN is not applicable to cyclic queries, thus some entries are marked “NA”. IVMϵ is not a concrete algorithm but rather a general methodology. It is not clear how to manually set it up for every query, thus some entries are marked “Open”. In query complexity, MVIVM has a query-dependant factor of $k!$, where $k$ is the number of input relations.
@@
In the following, we illustrate how the aforementioned IVM approaches maintain the 3-path query $Q_{3p}$ and the triangle query $Q_{\triangle}$.
Convention
To simplify the following presentation, we denote a single tuple update to some relation $R(A,B)$ by $\delta R(a,b)$, which expresses that the tuple $(a,b)$ is inserted to or deleted from $R$. Whenever it simplifies the explanation of an update procedure, we do not distinguish between an insert and a delete and describe the main computation steps common to both cases.
@@ A.4. IVMϵ
Updates can change the degrees of values, which means that a light value can become heavy and vice-versa. The original IVMϵ algorithm moves tuples between the heavy and the light part of a relation whenever the degrees of the values in the tuples pass certain thresholds (Kara et al., 2020a). This leads to additional time cost that can be amortized over update sequences such that the amortized update time remains $\mathcal{O}(|\mathcal{D}|^{1/2})$.
@@ A.6. MVIVM
Finally, we highlight some of the key insights of our approach, MVIVM, that enables it to maintain the triangle query $Q_{\triangle}$ in amortized time $\widetilde{\mathcal{O}}(|\mathcal{D}|^{1/2})$. Example 7.2 presents the high-level idea while Appendix E.1 explains the technical subtleties that arise. At a high-level, our approach maintains $Q_{\triangle}$ by constructing and maintaining the six multivariate extension components $\widehat{Q}_{123},\ldots,\widehat{Q}_{321}$ given by Eq. (8) and (39). Each of these component queries is maintained using Theorem 4.1 along with a variety of technical lemmas described in Appendix E.1. It can be mechanically verified that each of these components has a fractional hypertree width of $\frac{3}{2}$, leading to our amortized update time of $\widetilde{\mathcal{O}}(|\mathcal{D}|^{1/2})$. Let’s take the first component $\widehat{Q}_{123}$ as an example:
@@
Recall that $Z_{1}$, $Z_{2}$, and $Z_{3}$ are temporal attributes representing the lifespans of tuples in $R$, $S$, and $T$. In particular, the lifespan of a tuple in $T(A,C)$ is represented by the concatenation of $Z_{1}\circ Z_{2}\circ Z_{3}$; see Example 6.4. The above tree decomposition corresponds to a query plan where we are projecting $Z_{3}$ out of the child bag before joining with the root bag. Projecting $Z_{3}$ out of the child bag corresponds to “expanding” the lifespans of tuples of $T(A,C)$ from the lifespan $Z_{1}\circ Z_{2}\circ Z_{3}$ to the longer lifespan $Z_{1}\circ Z_{2}$. What this means algorithmically is that we are lazily processing deletes from $T$ in order to avoid doing excessive work for each delete, especially in the case where the same tuple from $T$ is inserted and deleted many times. In particular, a sequence of repeated insert-delete of the same tuple of $T$ corresponds to lifespans whose encodings may share the same prefix $Z_{1}\circ Z_{2}$. Hence, we don’t need to propagate changes from the child bag to the root bag for every such insert-delete because the lifespan $Z_{1}\circ Z_{2}$ remains in the child bag throughout all these insert-deletes. Without this lazy processing, each such repeated insert/delete into $T$ would have required linear update time, thus we would have failed to meet our target update time of $\widetilde{\mathcal{O}}(|\mathcal{D}|^{1/2})$. This implicit lazy processing is just one example of sophisticated behavior that comes automatically out of our approach.
@@ Appendix C Missing Details in Section 5
In this section, we explain in detail our results on IVM in the insert-only setting, where no deletes are allowed. The main goal is to prove Theorem 4.1 (as well as Proposition 4.3). The results in this section also serve as building blocks for our results on the insert-delete setting in Section E.
@@ C.1. Preliminaries: The Query Decomposition Lemma
Lemma 9 (Query Decomposition Lemma (Ngo et al., 2014; Ngo, 2018)).
@@
(27) $\sum_{\bm{y}\in\text{\sf Dom}(\bm{Y})}\underbrace{\prod_{R(\bm{X})\in\text{\sf at}(Q)}|R(\bm{X})\ltimes\bm{y}|^{\lambda_{R(\bm{X})}}}_{\text{AGM-bound of $Q\ltimes\bm{y}$}}\leq\underbrace{\prod_{R(\bm{X})\in\text{\sf at}(Q)}|R|^{\lambda_{R(\bm{X})}}}_{\text{AGM-bound of $Q$}}$